An intravenous solution of Dextrose 5% and 0.45% Normal Saline is infusing at a rate of 28 gtt/minute. The drop factor is 10 gtt/mL and there is 700 mL of fluid remaining in the bag. How long will it take for the fluid to infuse into the client?

Round to the nearest tenth decimal point.
Answer #1

Total time for infusion =total volume /flow rate

flow rate =28gtt/min

=28/10ml(drop factor is 10gtt/ml) =2.8ml

Therefore time for infusion =700/2.8=250 min